# Efficient Indirect LLM Jailbreak via Multimodal-LLM Jailbreak

### v2

@@ Abstract
This paper focuses on jailbreaking attacks against large language models (LLMs), eliciting them to generate objectionable content in response to harmful user queries. Unlike previous LLM-jailbreak methods that directly orient to LLMs, our approach begins by constructing a multimodal large language model (MLLM) built upon the target LLM. Subsequently, we perform an efficient MLLM jailbreak and obtain a jailbreaking embedding. Finally, we convert the embedding into a textual jailbreaking suffix to carry out the jailbreak of target LLM. Compared to the direct LLM-jailbreak methods, our indirect jailbreaking approach is more efficient, as MLLMs are more vulnerable to jailbreak than pure LLMs. Additionally, to improve the attack success rate of jailbreak, we propose an image-text semantic matching scheme to identify a suitable initial input. Extensive experiments demonstrate that our approach surpasses current state-of-the-art jailbreak methods in terms of both efficiency and effectiveness. Moreover, our approach exhibits superior cross-class generalization abilities. The code is available here.
1 Introduction
@@
Recently, large language models (LLMs) such as ChatGPT [7] have been widely deployed. These models exhibit advanced general abilities but also pose serious safety risks such as truthfulness, toxicity, and bias [17, 26, 34, 3, 9]. Typically, there exists a type of attack called jailbreaking attack, which can elicit LLMs to generate objectionable content in response to users’ harmful queries. For example, a pioneering work [45] has found that a specific prompt suffix allows the jailbreak of most popular LLMs. However, the efficiency of those methods is recognized to be relatively low, primarily attributed to the challenges of discrete optimization in finding the textual jailbreaking suffix (JBtxt).
@@
On the other hand, there is a surge of interest in multimodal large language models (MLLMs), which enable users to provide images as input to LLMs. [21, 14, 44, 11, 4, 43, 6, 37, 1]. Consequently, research on jailbreak has been extended from LLMs to MLLMs. Furthermore, it has been demonstrated that performing MLLM-jailbreak is easier and more efficient than performing LLM-jailbreak [31, 28]. This is largely because finding jailbreaking images within continuous pixel spaces is significantly easier and more flexible than identifying jailbreaking text within discrete token spaces.
@@
Inspired by that, we propose an efficient indirect LLM-jailbreaking approach that constructs an MLLM built upon the target LLM and subsequently performs MLLM-jailbreak. Specifically, we adopt a widely used LLM-jailbreaking strategy [45], which seeks to identify a specific prompt suffix, denoted as JBtxt. When appended to harmful queries, this JBtxt suffix can elicit LLMs to generate objectionable content. Our contribution lies in efficiently finding such JBtxt by leveraging the MLLM-jailbreak process. As shown in Fig.1, the workflow of our approach is as follows: (1) Given a target LLM to be jailbroke (e.g., LLaMA2), we first construct an MLLM by integrating a visual module into the target LLM. (2) We then perform an efficient MLLM-jailbreak. (3) Instead of using the jailbreaking image (referred to as JBimg), we obtain the output features from the visual module—namely, the Jailbreaking embeddings (JBemb)—and convert them into textual strings using our De-embedding and De-tokenization operations. (4) We regard these textual strings as JBtxt and append them to the harmful queries to carry out the jailbreak of the target LLM.
@@
Essentially, our approach leverages an MLLM-jailbreak to achieve the ultimate goal of LLM jailbreak. This indirect jailbreaking scheme offers flexibility for both white-box and black-box jailbreak. In the context of white-box jailbreak, the process of converting JBemb to JBtxt allows us to generate a set of candidate JBtxt. Unlike GCG [45], which produces only a single JBtxt, our approach outputs a diverse set of high-quality JBtxt, significantly enhancing the attack success rate (ASR) of jailbreak.
@@
Regarding the MLLM-jailbreak, we observe that the ASR of jailbreak is closely related to the initial input image (namely JBinit). By selecting an appropriate JBinit, we can significantly enhance the MLLM-jailbreaking ASR. To achieve this, we propose an image-text semantic matching scheme to identify this suitable JBinit. Our scheme attempts to align the embedding of JBinit with that of harmful queries. The underlying intuition is that aligned embeddings between the image and text inputs are more effective at triggering the model’s cross-attention mechanism, thereby exerting a stronger influence on the answer generation process. As a result, our JBinit can effectively shift the model’s response from “Sorry, I cannot" to “Sure, here is", thus achieving a successful jailbreak.
Our approach is also related to embedding-based LLM jailbreak methods [42], which focus on direct optimizing token embeddings. Compared to token-based methods such as GCG, these approaches are more efficient, as they rely on continuous optimization. However, previous embedding-based jailbreaks have been shown to be largely ineffective, primarily because the obtained embeddings often lack corresponding tokens [45]. In contrast, our method does not directly optimize JBemb. Instead, we indirectly obtain JBemb by optimizing JBimg. This procedure can be viewed as regularizing the optimization of JBemb through the use of the visual module. Since the visual module (e.g., a CLIP encoder) is trained with an image-text alignment objective, it increases the likelihood that our JBemb corresponds to valid tokens. Consequently, our approach surpasses traditional embedding-based jailbreak methods in terms of effectiveness.
@@
We conduct extensive experiments demonstrating that our approach outperforms current state-of-the-art jailbreaking methods in both efficiency and effectiveness. Efficiency is a key advantage of our approach. For instance, for the same jailbreaking tasks, our method requires only approximately 3% of GCG’s runtime. Furthermore, our approach provides notable flexibility for black-box jailbreak, enabling successful jailbreaks of recent LLMs such as Mistral-v0.3, Gemma2, Deepseek-R1, Grok-3, and DouBao1.5. Furthermore, it exhibits superior cross-class generalization ability of jailbreak. This finding suggests that improving the ASR for a specific class can benefit not only from data within the class itself, but also from data drawn from semantically related classes.
2 Related Work
LLM-jailbreak. Prior work on LLM jailbreak primarily focuses on crafting input prompts—referred to as jailbreaking prompts—to circumvent model alignment constraints. Early approaches relied on manually designing prompts with specific strategies, such as role-playing [26], prompt injection, and privilege escalation [33]. More recently, research has shifted toward automating the generation of jailbreak prompts to reduce the reliance on manual efforts.
Existing jailbreak techniques can be broadly categorized into two types. The first is token-based jailbreak, which directly optimizes discrete token sequences. For instance, [16] leverage time-based SQL injection techniques to automatically generate jailbreak prompts. [22] introduce autoDAN, a hierarchical genetic algorithm that evolves stealthy and effective prompts. The GCG method proposed by [45] appends specific suffixes to input queries to enable automated jailbreaks.
The second category is embedding-based jailbreak, which starts by optimizing token embeddings and subsequently projects them back into discrete tokens. A notable challenge with this approach is that the optimized embeddings often do not correspond to any valid token. To address this, Prompts Made Easy (PEZ) [42] adopts a quantized optimization strategy that adjusts continuous embeddings while ensuring they can be effectively mapped back to tokens.
MLLM-jailbreak. Recently, the safety risks posed by multimodal large language models (MLLMs) have attracted growing concern. Several jailbreak techniques specifically designed for MLLMs have emerged. For instance, a pioneering study by Carlini [10] demonstrated that adversarial prompts in the form of continuous-domain images can effectively induce MLLMs to generate harmful or toxic content. Similarly, [32] proposed decomposing a malicious prompt into a benign textual query combined with malicious triggers—either OCR-based textual triggers or visual triggers—exploiting the joint embedding space to enable the jailbreak. These methods collectively reveal that MLLM jailbreaking achieves significantly higher success rates, often surpassing those of traditional LLM jailbreak strategies.
3 Our Approach
@@ 3.2 MLLM Construction
3.3 MLLM Jailbreak
We adopt a strategy for MLLM-jailbreak analogous to that used in LLM-jailbreak, i.e., encouraging the MLLM to generate a response that begins with a positive affirmation such as “Sure, here is a (content of query)." Our method is inspired by adversarial attacks, which typically perturb visual inputs to manipulate a model’s output. However, jailbreak fundamentally differs from adversarial attacks in that it targets a generative task rather than a discriminative one. To bridge this gap, we adapt adversarial techniques by replacing the conventional cross-entropy loss with a maximum likelihood objective, thereby encouraging the MLLM to produce affirmative responses. This objective directly maximizes the likelihood of the desired output, making it better suited for the generative nature of jailbreak.
@@ 3.3 MLLM Jailbreak
$\mathop{\max}_{\delta}\sum_{i=0}^{M}log(p(a_{i}|q_{i},\widetilde{x}))$ (1)
@@
$\qquad||\delta||_{p}<\epsilon$
where $p(a_{i}|q_{i},\widetilde{x})$ represents the likelihood of the MLLM generating the answer $a_{i}$ when provided with the image $\widetilde{x}$ and the text query $q_{i}$. Here, $\epsilon$ denotes the attack budget, and $\widetilde{x}$ refers to the jailbreaking image (JBimg). The optimization process is performed over a set of $M$ query-answer pairs $\{(q_{i},a_{i}),i=0,...,M\}$. This problem can be effectively solved by adapting the Projected Gradient Descent (PGD) algorithm [24].
3.4 LLM Jailbreak
@@
For each embedding vector $e_{l}$ in the sequence $(e_{1},...,e_{L})$, we identify the top-$K$ closest embedding vectors $\hat{e}^{k}_{l}$, where $k=1,...,K$. Performing this for all $e_{l}$ results in a $K\times L$ embedding pool $\{\hat{e}_{l}^{k}\}_{k=1,l=1}^{K,L}$ and a corresponding $K\times L$ token pool $\{\hat{t}_{l}^{k}\}_{k=1,l=1}^{K,L}$. The De-tokenization operation then converts these tokens into natural language words, yielding a $K\times L$ word pool $\{\hat{w}_{l}^{k}\}_{k=1,l=1}^{K,L}$. Finally, we randomly sample several word sequences from this word pool—each sampled sequence constituting a candidate JBtxt.
@@
3.5 Finding an Appropriate JBInit
We observe that initializing the MLLM-jailbreak algorithm with a suitable starting point, namely JBInit, can substantially enhance the ASR of the MLLM-jailbreak. To this end, we propose an image-text semantic matching scheme. Specifically, our goal is to ensure that the embedding of JBInit is close to the embedding of the harmful queries. The rationale behind this scheme is that, for MLLM-jailbreak to succeed, the JBInit must exert a substantial influence on the LLM’s answer generation process. Particularly, it needs to turn a typical refusal response like “Sorry, I cannot" into a positive affirmation such as “Sure, here is". By ensuring that the embedding of JBInit closely aligns with that of the query, we can effectively trigger the model’s cross-attention mechanism, thereby exerting a stronger influence on the answer generation process.
@@ 3.5 Finding an Appropriate JBInit
4 Evaluation
Datasets. Several datasets have been introduced for evaluating jailbreak attacks, such as AdvBench [45], which includes a diverse range of harmful behaviors, including violence, financial crimes, drug-related crimes, and more. However, existing evaluation protocols often aggregate all these behaviors into a single pool during assessment.
@@ 4 Evaluation
Specifically, we categorize AdvBench dataset into nine classes: “unlawful violence", “financial crimes", “property crimes", “drug crimes", “weapons crimes", “cyber crimes", “hate", “Suicide and Self-Harm", and “Fake info". Furthermore, since our approach involves MLLM-jailbreak, we extend the AdvBench dataset into a multimodal version, termed AdvBench-M, by augmenting it with images. Specifically, for each harmful class, we retrieve 20 semantically relevant images from the Internet using the Google search engine. We then employ CLIP ViT-L/14 [30] to select the top 10 images that best align with the queries associated with each harmful class.
In addition to AdvBench, we conducted extensive experiments on HarmBench dataset [25]. HarmBench consists of 400 harmful behavior instances. Notably, HarmBench encompasses several distinct behavior categories not included in AdvBench, such as copyright violations. To ensure a rigorous evaluation of our proposed method, we strictly followed the evaluation protocols outlined in the HarmBench paper.
Test Models. We evaluate our approach with both white-box and black-box jailbreaking settings. For the white-box setting, we conduct evaluations on open-source models, including LLaMA2-Chat-7B [40], InternLM2.5-7B [8], and Qwen2.5-3B-Instruct [39]. Specifically, for LLaMA2-Chat-7B, we utilize MiniGPT-4 [44] as the corresponding MLLM; for InternLM2.5-7B, we utilize InternVL2.0-8B [13] as the corresponding MLLM; and for Qwen2.5-3B-Instruct, we utilize InternVL2.5-4B [12] as its corresponding MLLM.
For the black-box setting, we generate the JBtxt by using InternVL2.0 as MLLM, and subsequently employ these textual suffixes to jailbreak closed-source models, including Mistral-v0.3 [19], Gemma2 [36], ChatGLM4 [18], Deepseek-R1 [15], Grok-3 [2], and DouBao1.5 [35].
Assessment. Evaluating the success of jailbreak is inherently challenging. The objective of jailbreak is to elicit the LLM to produce any response that complies with a harmful query, making it difficult to define a specific ground truth for each query. Moreover, we have observed cases where responses begin with seemingly affirmative phrases (e.g., “Sure, here is a…") but ultimately decline to fulfill the query. Consequently, prior methods often rely on manual inspection to judge jailbreak success. To overcome this limitation, we propose using the LLaMA Guard 2 tool [38] to automatically assess the success of jailbreak attempts.
Table 1: White-box jailbreak of LLaMA2 (via MiniGPT4-Jailbreak) on AdvBench-M dataset. Performance is evaluated using ASR (%).
@@
4.1 White-box Jailbreak
We compare our approach with both token-based jailbreaking methods (i.e., AutoPrompt[27], AutoDAN [23], GCG [45], AdvAttack [5]) and embedding-based methods (i.e., PEZ [42], Soft Prompting [29]). Generally, the first category exhibits strong performance. GCG directly optimizes the JBtxt, while AutoDAN extends GCG by enhancing the readability of the generated JBtxt. AdvAttack introduces an adaptive strategy combining “Prompt + Random Search (RS) + Self-transfer," where the self-transfer component is employed to further boost jailbreaking effectiveness.
@@ 4.1 White-box Jailbreak
4.1.1 Compare to Token-based Jailbreak
@@
4.1.2 Compare to Embedding-based Jailbreak
Compared to token-based jailbreak, embedding-based jailbreaking methods can leverage straightforward and efficient continuous optimization, since token embeddings are continuous variables. This technique is often referred to as “soft prompting"[20, 29, 41] in other literature and has demonstrated advantages in the Prompt Engineering domain. However, as pointed out in [45], this type of method is ineffective in LLM-jailbreak because the obtained embeddings (i.e., soft prompts) often have no corresponding token. In our experiments, we reached a similar conclusion to [45], where the PEZ method completely failed to achieve any successful jailbreaks, i.e., $\text{ASR}\approx 0$ for all classes. Thus, the results of PEZ are omitted from Tables 1 and 2. The Soft Prompting (SoftP) method [29] achieves slightly better performance than PEZ; however, its ASR remains lower than that of token-based methods, as shown in Tables 1 and 2.
@@ 4.1.2 Compare to Embedding-based Jailbreak
4.1.3 Comparison on HarmBench
Besides the AdvBench-M dataset, we also evaluate jailbreaking performance on the HarmBench dataset [25]. From Table 3, our approach significantly outperforms all baseline methods. Compared to AdvAttack, which shows strong performance on HarmBench, our method achieves a 4.5% improvement on LLaMA-2 (72.0% vs. 67.5% ASR on the entire dataset) and a 1.8% improvement on InternLM-2.5 (63.0% vs. 61.2% ASR), fully demonstrating the effectiveness of our approach.
Class Black-box
Mistral-v0.3 Gemma2 Deepseek-R1 Grok-3 DouBao1.5
Unlawful violence 72.2 88.9 72.2 38.9 5.6
Financial crimes 96.4 92.8 89.2 62.5 39.2
Property crimes 93.7 92.1 78.1 67.2 31.2
Drug crimes 86.7 80.0 86.7 26.7 6.7
Weapons crimes 88.9 44.4 77.8 5.57 11.1
Cyber crimes 87.5 85.0 87.5 77.5 35.0
Hate 66.7 53.3 40.0 60.0 13.3
Suicide/Self-Harm 22.2 61.1 16.7 66.7 0.0
Fake info 80.0 93.3 86.7 33.3 40.0
Average ASR 77.1 76.7 70.5 48.7 20.1
4.2 Black-box Jailbreak
@@
4.3 Discussion
4.3.1 Image-text Semantic Matching
In our approach, MLLM-jailbreak is achieved by perturbing the JBinit with a perturbation $\Delta$, i.e., $\text{JBimg}=\text{JBinit}+\Delta$. In this section, we illustrate that finding an appropriate JBinit is crucial for both successful MLLM-jailbreak and subsequent LLM-jailbreak.
In our experiments, we compare three distinct schemes for finding JBInit. The first scheme involves randomly sampling one image from all images in the AdvBench-M dataset, regardless of its harmful class. The second scheme restricts the sampling to images within the same harmful class as the query, and further employs CLIP ViT-L/14 to identify the image that best aligns with the semantics of the corresponding class. We refer to this as the ranking-based scheme. The third scheme is our proposed image-text semantic matching approach.

### v3

# Efficient LLM-Jailbreaking via Multimodal-LLM Jailbreak

@@ Abstract
This paper focuses on jailbreaking attacks against large language models (LLMs), eliciting them to generate objectionable content in response to harmful user queries. Unlike previous LLM-jailbreak methods that directly orient to LLMs, our approach begins by constructing a multimodal large language model (MLLM) built upon the target LLM. Subsequently, we perform an efficient MLLM jailbreak and obtain a jailbreaking embedding. Finally, we convert the embedding into a textual jailbreaking suffix to carry out the jailbreak of target LLM. Compared to the direct LLM-jailbreak methods, our indirect jailbreaking approach is more efficient, as MLLMs are more vulnerable to jailbreak than pure LLMs. Additionally, to improve the attack success rate of jailbreak, we propose an image-text semantic matching scheme to identify a suitable initial input. Extensive experiments demonstrate that our approach surpasses current state-of-the-art jailbreak methods in terms of both efficiency and effectiveness. Moreover, our approach exhibits superior cross-class generalization abilities.
Codehttps://github.com/nobody235/LLM-jailbreak
Introduction
@@
Recently, large language models (LLMs) such as ChatGPT (brown2020language) have been widely deployed. These models exhibit advanced general abilities but also pose serious safety risks such as truthfulness, toxicity, and bias (gehman2020realtoxicityprompts; perez2022red; sheng2019woman; abid2021persistent; carlini2021extracting). Typically, there exists a type of attack called jailbreaking attack, which can elicit LLMs to generate objectionable content in response to users’ harmful queries. For example, a pioneering work (zou2023universal) has found that a specific prompt suffix allows the jailbreak of most popular LLMs. However, the efficiency of those methods is recognized to be relatively low, primarily attributed to the challenges of discrete optimization in finding the textual jailbreaking suffix (JBtxt).
@@
On the other hand, there is a surge of interest in multimodal large language models (MLLMs), which enable users to provide images as input to LLMs. (liu2024visual; dai2024instructblip; zhu2023minigpt; chen2023minigpt; alayrac2022flamingo; ye2023mplug; bai2023qwen; team2023gemini; 2023GPT4VisionSC). Consequently, research on jailbreak has been extended from LLMs to MLLMs. Furthermore, it has been demonstrated that performing MLLM-jailbreak is easier and more efficient than performing LLM-jailbreak (shayegani2023jailbreak; qi2024visual). This is largely because finding jailbreaking images within continuous pixel spaces is significantly easier and more flexible than identifying jailbreaking text within discrete token spaces.
@@
Inspired by that, we propose an efficient indirect LLM-jailbreaking approach that constructs an MLLM built upon the target LLM and subsequently performs MLLM-jailbreak. Specifically, we adopt a widely used LLM-jailbreaking strategy (zou2023universal), which seeks to identify a specific prompt suffix, denoted as JBtxt. When appended to harmful queries, this JBtxt suffix can elicit LLMs to generate objectionable content. Our contribution lies in efficiently finding such JBtxt by leveraging the MLLM-jailbreak process. As shown in Fig.1, the workflow of our approach is as follows: (1) Given a target LLM to be jailbroke (e.g., LLaMA2), we first construct an MLLM by integrating a visual module into the target LLM. (2) We then perform an efficient MLLM-jailbreak. (3) Instead of using the jailbreaking image (referred to as JBimg), we obtain the output features from the visual module—namely, the Jailbreaking embeddings (JBemb)—and convert them into textual strings using our De-embedding and De-tokenization operations. (4) We regard these textual strings as JBtxt and append them to the harmful queries to carry out the jailbreak of the target LLM.
@@
Essentially, our approach leverages an MLLM-jailbreak to achieve the ultimate goal of LLM jailbreak. This indirect jailbreaking scheme offers flexibility for both white-box and black-box jailbreak. In the context of white-box jailbreak, the process of converting JBemb to JBtxt allows us to generate a set of candidate JBtxt. Unlike GCG (zou2023universal), which produces only a single JBtxt, our approach outputs a diverse set of high-quality JBtxt, significantly enhancing the attack success rate (ASR) of jailbreak.
@@
Regarding the MLLM-jailbreak, we observe that the ASR of jailbreak is closely related to the initial input image (namely JBinit). By selecting an appropriate JBinit, we can significantly enhance the MLLM-jailbreaking ASR. To achieve this, we propose an image-text semantic matching scheme to identify this suitable JBinit. Our scheme attempts to align the embedding of JBinit with that of harmful queries. The underlying intuition is that aligned embeddings between the image and text inputs are more effective at triggering the model’s cross-attention mechanism, thereby exerting a stronger influence on the answer generation process. As a result, our JBinit can effectively shift the model’s response from “Sorry, I cannot” to “Sure, here is”, thus achieving a successful jailbreak.
Our approach is also related to embedding-based LLM jailbreak methods (wen2024hard), which focus on direct optimizing token embeddings. Compared to token-based methods such as GCG, these approaches are more efficient, as they rely on continuous optimization. However, previous embedding-based jailbreaks have been shown to be largely ineffective, primarily because the obtained embeddings often lack corresponding tokens (zou2023universal). In contrast, our method does not directly optimize JBemb. Instead, we indirectly obtain JBemb by optimizing JBimg. This procedure can be viewed as regularizing the optimization of JBemb through the use of the visual module. Since the visual module (e.g., a CLIP encoder) is trained with an image-text alignment objective, it increases the likelihood that our JBemb corresponds to valid tokens. Consequently, our approach surpasses traditional embedding-based jailbreak methods in terms of effectiveness.
@@
We conduct extensive experiments demonstrating that our approach outperforms current state-of-the-art jailbreaking methods in both efficiency and effectiveness. Efficiency is a key advantage of our approach. For instance, for the same jailbreaking tasks, our method requires only approximately 3% of GCG’s runtime. Furthermore, our approach provides notable flexibility for black-box jailbreak, enabling successful jailbreaks of recent LLMs such as Mistral-v0.3, Gemma2, Deepseek-R1, Grok-3, and DouBao1.5. Furthermore, it exhibits superior cross-class generalization ability of jailbreak. This finding suggests that improving the ASR for a specific class can benefit not only from data within the class itself, but also from data drawn from related classes.
Our Approach
@@
MLLM Jailbreak
We adopt a strategy for MLLM-jailbreak analogous to that used in LLM-jailbreak, i.e., encouraging the MLLM to generate a response that begins with a positive affirmation such as “Sure, here is a (content of query).” Our method is inspired by adversarial attacks, which typically perturb visual inputs to manipulate a model’s output. However, jailbreak fundamentally differs from adversarial attacks in that it targets a generative task rather than a discriminative one. To bridge this gap, we adapt adversarial techniques by replacing the conventional cross-entropy loss with a maximum likelihood objective, thereby encouraging the MLLM to produce affirmative responses. This objective directly maximizes the likelihood of the desired output, making it better suited for the generative nature of jailbreak.
@@
$)\displaystyle\mathop{\max}_{\delta}\sum_{i=0}^{M}log(p(a_{i}|q_{i},\widetilde{x}$)) (1)
$\text{s.t. }\widetilde{x}\in[0,255]^{d},\widetilde{x}=x_{init}+\delta$
@@
[…]
LLM Jailbreak
@@
For each embedding vector $e_{l}$ in the sequence $(e_{1},...,e_{L})$, we identify the top-$K$ closest embedding vectors $\hat{e}^{k}_{l}$, where k=1,,Kk=1,...,K. Performing this for all $e_{l}$ results in a $K\times L$ embedding pool $\{\hat{e}_{l}^{k}\}_{k=1,l=1}^{K,L}$ and a corresponding $K\times L$ token pool $\{\hat{t}_{l}^{k}\}_{k=1,l=1}^{K,L}$. The De-tokenization operation then converts these tokens into natural language words, yielding a $K\times L$ word pool $\{\hat{w}_{l}^{k}\}_{k=1,l=1}^{K,L}$. Finally, we randomly sample several word sequences from this word pool—each sampled sequence constituting a candidate JBtxt.
@@
Finding an Appropriate JBInit
We observe that initializing the MLLM-jailbreak algorithm with a suitable starting point, namely JBInit, can substantially enhance the ASR of the MLLM-jailbreak. To this end, we propose an image-text semantic matching scheme. Specifically, our goal is to ensure that the embedding of JBInit is close to the embedding of the harmful queries. The rationale behind this scheme is that, for MLLM-jailbreak to succeed, the JBInit must exert a substantial influence on the LLM’s answer generation process. Particularly, it needs to turn a typical refusal response like “Sorry, I cannot” into a positive affirmation such as “Sure, here is”. By ensuring that the embedding of JBInit closely aligns with that of the query, we can effectively trigger the model’s cross-attention mechanism, thereby exerting a stronger influence on the answer generation process.
@@
Evaluation
Datasets. Several datasets have been introduced for evaluating jailbreak attacks, such as AdvBench (zou2023universal), which includes a diverse range of harmful behaviors, including violence, financial crimes, drug-related crimes, and more. However, existing evaluation protocols often aggregate all these behaviors into a single pool during assessment.
@@
Specifically, we categorize AdvBench dataset into nine classes: “Unlawful Violence”, “Financial Crimes”, “Property Crimes”, “Drug Crimes”, “Weapons Crimes”, “Cyber Crimes”, “Hate”, “Suicide and Self-Harm”, and “Fake Info”. Furthermore, since our approach involves MLLM-jailbreak, we extend the AdvBench dataset into a multimodal version, termed AdvBench-M, by augmenting it with images. Specifically, for each harmful class, we retrieve 20 semantically relevant images from the Internet using the Google search engine. We then employ CLIP ViT-L/14 (radford2021learning) to select the top 10 images that best align with the queries associated with each harmful class.
@@
Table 1: White-box jailbreak of LLaMA2 (via MiniGPT4-Jailbreak) on AdvBench-M. Performance is evaluated using ASR.
@@
In addition to AdvBench, we conducted extensive experiments on HarmBench dataset (HarmBench). HarmBench consists of 400 harmful behavior instances. Notably, HarmBench encompasses several distinct behavior categories not included in AdvBench, such as “Copyright Violations”. To ensure a rigorous evaluation of our proposed method, we strictly followed the evaluation protocols outlined in the HarmBench paper.
Test Models. We evaluate our approach with both white-box and black-box jailbreaking settings. For the white-box setting, we conduct evaluations on open-source models, including LLaMA2-Chat-7B (touvron2023llama), InternLM2.5-7B (cai2024internlm2), and Qwen2.5-3B-Instruct (qwen2025qwen25). Specifically, for LLaMA2-Chat-7B, we utilize MiniGPT-4 (zhu2023minigpt) as the corresponding MLLM; for InternLM2.5-7B, we utilize InternVL2.0-8B (chen2024internvl) as the corresponding MLLM; and for Qwen2.5-3B-Instruct, we utilize InternVL2.5-4B (chen2024VL25) as its corresponding MLLM.
For the black-box setting, we generate the JBtxt by using InternVL2.0 as MLLM, and subsequently employ these textual suffixes to jailbreak closed-source models, including Mistral-v0.3 (jiang2023mistral), Gemma2 (gemmateam2024gemma2), ChatGLM4 (glm2024chatglm4), Deepseek-R1 (deepseekai2025deepseekr1), Grok-3 (Grok3), and DouBao1.5 (guo2025seed15vl).
Assessment. Evaluating the success of jailbreak is inherently challenging. The objective of jailbreak is to elicit the LLM to produce any response that complies with a harmful query, making it difficult to define a specific ground truth for each query. Moreover, we have observed cases where responses begin with seemingly affirmative phrases (e.g., “Sure, here is a…”) but ultimately decline to fulfill the query. Consequently, prior methods often rely on manual inspection to judge jailbreak success. To overcome this limitation, we propose using the LLaMA Guard 2 tool (metallamaguard2) to automatically assess the success of jailbreak attempts.
White-box Jailbreak
We compare our approach with both token-based jailbreaking methods (i.e., AutoPrompt(pryzant2023automatic), AutoDAN (liu2024AutoDAN), GCG (zou2023universal), AdvAttack (andriushchenko2024PRS)) and embedding-based methods (i.e., PEZ (wen2024hard), Soft Prompting (qin2021learning)). Generally, the first category exhibits strong performance. GCG directly optimizes the JBtxt, while AutoDAN extends GCG by enhancing the readability of the generated JBtxt. AdvAttack introduces an adaptive strategy combining “Prompt + Random Search (RS) + Self-transfer,” where the self-transfer component is employed to further boost jailbreaking effectiveness.
@@
Compare to Token-based Jailbreak
@@
Compare to Embedding-based Jailbreak
Compared to token-based jailbreak, embedding-based jailbreaking methods can leverage straightforward and efficient continuous optimization, since token embeddings are continuous variables. This technique is often referred to as “soft prompting”(lester2021power; qin2021learning; vu2021spot) in other literature and has demonstrated advantages in the Prompt Engineering domain. However, as pointed out in (zou2023universal), this type of method is ineffective in LLM-jailbreak because the obtained embeddings (i.e., soft prompts) often have no corresponding token. In our experiments, we reached a similar conclusion to (zou2023universal), where the PEZ method completely failed to achieve any successful jailbreaks, i.e., $\text{ASR}\approx 0$ for all classes. Thus, the results of PEZ are omitted from Tables 1 and 2. The Soft Prompting (SoftP) method (qin2021learning) achieves slightly better performance than PEZ; however, its ASR remains lower than that of token-based methods, as shown in Tables 1 and 2.
@@
Comparison on HarmBench
Besides the AdvBench-M dataset, we also evaluate jailbreaking performance on the HarmBench dataset (HarmBench). From Table 3, our approach significantly outperforms all baseline methods. Compared to AdvAttack, which shows strong performance on HarmBench, our method achieves a 4.5% improvement on LLaMA2 (72.0% vs. 67.5% ASR on the entire dataset) and a 1.8% improvement on InternLM2.5 (63.0% vs. 61.2% ASR), fully demonstrating the effectiveness of our approach.
@@
Class Black-box
Mistral Gemma2 Deepseek-R1 Grok-3 DouBao1.5
Violence 72.2 88.9 72.2 38.9 5.6
Financial 96.4 92.8 89.2 62.5 39.2
Property 93.7 92.1 78.1 67.2 31.2
Drug 86.7 80.0 86.7 26.7 6.7
Weapons 88.9 44.4 77.8 5.57 11.1
Cyber 87.5 85.0 87.5 77.5 35.0
Hate 66.7 53.3 40.0 60.0 13.3
Self-Harm 22.2 61.1 16.7 66.7 0.0
Fake info 80.0 93.3 86.7 33.3 40.0
Avg ASR 77.1 76.7 70.5 48.7 20.1
Black-box Jailbreak
Discussion
Image-text Semantic Matching
In our approach, MLLM-jailbreak is achieved by perturbing the JBinit with a perturbation $\Delta$, i.e., $\text{JBimg}=\text{JBinit}+\Delta$. In this section, we illustrate that finding an appropriate JBinit is crucial for both successful MLLM-jailbreak and subsequent LLM-jailbreak. In our experiments, we compare three distinct schemes for finding JBInit. The first scheme involves randomly sampling one image from all images in the AdvBench-M dataset, regardless of its harmful class. The second scheme restricts the sampling to images within the same harmful class as the query, and further employs CLIP ViT-L/14 to identify the image that best aligns with the semantics of the corresponding class. We refer to this as the ranking-based scheme. The third scheme is our proposed image-text semantic matching approach.